OpenStudy (anonymous):

Locate the absolute maximum value of the functionon the interval [0, 5].

OpenStudy (anonymous):

check if any maxima lies in the interval [0,5] by finding f'(x)=0...after getting maximas and minimas check if f''(x)<0 then its maxima then just put that value of x in the f(x)

OpenStudy (anonymous):

f'(x)= 3x^2 -27

OpenStudy (anonymous):

when does \(3x^2-27=3(x^2-9)=3(x+3)(x-3)=0\)?

OpenStudy (anonymous):

When x=3 and x=-3

OpenStudy (anonymous):

Is there even a maximum value of this function?

OpenStudy (anonymous):

to find absolute maximum, you need to check at x = 2, x = -3 and boundary points

OpenStudy (anonymous):

*x=3

OpenStudy (anonymous):

okay so would my answer be k-54?

OpenStudy (anonymous):

That is what I got from f(x) at 3

OpenStudy (anonymous):

you need to find all the 4 values : f(3) f(-3) f(0) f(5) and pick the max value

OpenStudy (anonymous):

f(3)= k-54 f(-3)= k+54 f(0)=k f(5)= 206+k

OpenStudy (anonymous):

oh it is k at 0

OpenStudy (anonymous):

f(3) : k-54 f(-3) : k - 108 f(0) : k f(5) : k-10

OpenStudy (anonymous):

Okay. It must be f(0)

OpenStudy (anonymous):

absolute max = max [f(3), f(-3), f(0), f(5)] = k yes